OpenStudy (anonymous):

During a clothing store’s Bargain Days, the regular price for T-shirts is discounted by $5. There is a state sales tax of 5%, and the $5 discount is applied before the sales tax is calculated. a. Write an expression that shows the regular price r of a T-shirt minus the $5 discount. b. Write a rule for the function p(r) that expresses the final price p of a T-shirt with the discount applied and sales tax added. c. How much would you pay during Bargain Days for a shirt regularly priced at $15.50?

OpenStudy (anonymous):

For the first part I got r(0.05)

OpenStudy (anonymous):

Well to answer c you would pay $14.725 :p, or 14.73

OpenStudy (anonymous):

sorry, I did 5% of 15.50.

OpenStudy (anonymous):

You would actually pay 11.02 in order to answer c.

OpenStudy (mayankdevnani):

a. r – 5 b. p(r) = 1.05(r − 5) OR p(r) = r − 5 + 0.05(r − 5) c. $11.03

OpenStudy (mayankdevnani):

am i right @jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

yes !